OpenStudy (vlery):

In Problems 1 and 2, use a graphing utility to approximate the solutions of each equation rounded to two decimal places. All solutions lie between -10 and 10. 1. -x^3+3x+1=0 2. -x^4+7= x^2-2

OpenStudy (anonymous):

yes

OpenStudy (vlery):

no they need to be answered @KlOwNlOvE help me out I posted in the wrong section lol

OpenStudy (anonymous):

for 2. subtract x^2-2 from both sides -x^4+7-(x^2-2) distribute the negative -x^4-x^2+9 rewrite the equation in u=x^2 u^2=x^4 -u^2-u+9=0 divide both sides by -1 u^2+u-9=0 add 9 to both sides u^2+u=9 add (1/2)^2 to both sides u^2+u+1/4=37/4 complete the left hand side to square using the identity x^2+2ax+a^2=(x+a)^2 \[(u+\frac{ 1 }{ 4 })^{2}=\frac{ 37 }{ 4 }\] \[u+\frac{ 1 }{ 4 }=\sqrt{\frac{ 37 }{ 4 }}\] now subtract the 1/4 from both sides \[u=\frac{ \sqrt{37}-1 }{ 2 }\]

OpenStudy (anonymous):

@SolomonZelman seem right?